OpenStudy (anonymous):

How does 1/4 plus m = 1

OpenStudy (kobeni-chan):

m could equal 3/4, 1/4 + 3/4 = 1

jhonyy9 (jhonyy9):

1 -- + m = 1 4 1+4m=4 4m=3 m= 3/4

OpenStudy (anonymous):

m = 3/4 :D